OpenStudy (anonymous):

When 40 is added to the number of miles Karen ran last week, the result is the same as adding 10 to 4 times the number of miles she ran last week. How many miles did Karen run last week? mi____

OpenStudy (anonymous):

@SamsungFanBoy @AdamK

OpenStudy (adamk):

The equation would be 4x + 10 = x + 40

OpenStudy (adamk):

So solve that. Subtract 10 and you get 4x = x + 30 Then subtract x and you get 3x = 30 Divide by 3

OpenStudy (anonymous):

x=10 ? @AdamK

OpenStudy (adamk):

Yep! @edennoyes

OpenStudy (samsungfanboy):

Nice!